OpenStudy (anonymous):

HELP!I'm really bad at word problems. :This is about rate problems: Four years ago, the water in the apartment complex pool was emptied using a single pump. That took 33 hours. To get the job done more quickly this year, the apartment manager is going to use an additional pump that would take 22 hours to empty the entire pool on its own. If the pool contains 99,000 gallons of water, how much faster will the pool empty using both pumps at the same time than it did four years ago? To the nearest tenth of an hour, the pool will be emptied _____ hours faster.

OpenStudy (phi):

these use rate * time = amount we need to figure out the rates. pool contains 99,000 gallons single pump. That took 33 hours that means the rate for this pump is 99000 gals/33 hours which can be simplified to 3000 gal/hr (3000 gallons per hour) any idea what the rate of the second pump is ?

OpenStudy (anonymous):

um 99000/22?

OpenStudy (phi):

yes, and you should keep the units i.e. gals/hour and of course, simplify it

OpenStudy (anonymous):

Okay so its 4500 gals/ per hour ?

OpenStudy (phi):

yes. these pumps can operate in parallel (at the same time) so what is the combined rate of gals/ hour when both are working together ?

OpenStudy (anonymous):

so its 3000gals+4500gals

OpenStudy (phi):

which is 7500 gal/hour now use rate * time = 99,000 gals

OpenStudy (phi):

you know the rate, you know the amount of water you have to pump. You can solve for time

OpenStudy (anonymous):

so 99000/7500= 13.2

OpenStudy (anonymous):

Okay thanks! before you go can you help me with another question?

OpenStudy (phi):

make that \[ \frac{ 99000\ gal}{7500 \frac{ gal}{hour} }= 13.2 \ hours \]